hello,
could someone help me please, I have this questions..
1)consider a set of 900 of data values with a mean of 24 and standard deviation of 4.0
a)how many values are larger than 27.00
(report answer to nearest integer)
b) how many values are btween 23.00 and 28.00
(report answer to nearest integer)

Answer by stanbon(75887)

1)consider a set of 900 of data values with a mean of 24 and standard deviation of 4.0
a)how many values are larger than 27.00
z(27) = (27-24)/4 = 3/4
P(x > 27) = P(z > 3/4) = 0.2266
----
Answer: 0.2266*900 = 204
----------------------------------------
b) how many values are btween 23.00 and 28.00
z(28) = (28-24) = 4/4 = 1
P(27 < x < 28) = P(3/4 < z < 1) = normalcdf(3/4 , 1) = 0.06
------
Ans: 0.06*900 = 54
===================
Cheers,
Stan H.
===================